ten nn.hrend nbad tr T es went t, r t w n rws ther w.tthi em wh thei, ar ..ast ion.Os lat on th.
Answer #1

answer) its given the earth rotates 360o in 24 hr

so in 1 hour the earth rotates=360/24=15o

so in 10 hr=15*10=150o

so the answer is 150 degrees.